# Agile Tradespace Exploration for Space Rendezvous Mission Design via Transformers

### v1

@@ 7 Conclusion
This paper presents a rapid mission design framework for multi-objective spacecraft trajectory optimization under complex constraints. A Transformer model is trained to predict trajectories conditioned on flight time and target orbital elements, allowing it to generate diverse, high-quality solutions across various mission scenarios. The trained model is validated against real orbital debris data spanning a wide range of orbital elements, demonstrating strong performance as an initial guess for chance-constrained, passively safe rendezvous trajectory optimization, suitable for subsequent refinement. Moreover, results show that the generated trajectories themselves can act as an effective surrogate model of the mission profile, particularly in challenging cases where convex relaxations severely violate nonconvex constraints. By exploiting batched trajectory generation, the process achieves speeds approaching those of solving the convex relaxation itself. Overall, the findings highlight the potential of AI-based trajectory generation as a powerful tool for agile trade-space exploration in preliminary mission design. The proposed framework is readily extensible to a broad range of applications, from six-degree-of-freedom rendezvous to interplanetary transfers. Future research will focus on mitigating residual terminal constraint violations arising from autoregressive generation, for example by incorporating multiple-shooting-based trajectory inference.
@@ 7.2 SCP Parameters
Table 3: Hyperparaameters of $\texttt{SCVx}^{*}$.
Hyperparaemter Value

### v2

@@ 7 Conclusion
This paper presents a rapid mission design framework supported by the multi-objective spacecraft trajectory generation under complex nonconvex constraints. A key enabler is the transformer model that is trained to predict trajectories conditioned on flight time and target orbital elements, allowing it to generate diverse, high-quality solutions across various mission scenarios. The trained model is validated against real orbital debris data spanning a wide range of orbital elements, demonstrating strong performance as an initial guess for chance-constrained, passively safe rendezvous trajectory optimization, suitable for subsequent refinement. Moreover, results show that the generated trajectories themselves can act as an effective surrogate model of the mission profile, particularly in challenging cases where convex relaxations severely violate nonconvex constraints. By exploiting batched trajectory generation, the process achieves speeds approaching those of solving the convex relaxation itself. Overall, the findings highlight the potential of AI-based trajectory generation as a powerful tool for agile trade-space exploration in preliminary mission design. The proposed framework is readily extensible to a broad range of applications, from six-degree-of-freedom rendezvous to interplanetary transfers. Future research will focus on mitigating residual terminal constraint violations arising from autoregressive generation, for example by incorporating multiple-shooting-based trajectory inference. In addition, further reductions in inference runtime would enable onboard mission design and trade-space analysis, potentially allowing these capabilities to become an integral component of future spacecraft autonomy.
@@ 7.2 SCP Parameters
Table 3: Hyperparameters of $\texttt{SCVx}^{*}$.
Hyperparameter Value